Someone decides to replace six 60 watt filament bulbs with six 15 watt low energy bulbs . If energy costs 10 cu per unit and the bulbs are used , on average, for 4 hours per day what will the annual saving be?



1
Expert's answer
2021-06-14T14:56:44-0400

Gives

Bulb =6

Initial pawer

"P=\frac{W}{t}"

"E=P\times t"

"E=60\times4" watt "\times hour"

Number of bulbs=6


"E=6\times60\times4=1440watthour"


Monthly "E=30\times1440 =43200watthour"

Annual

"E=43200\times12 =518400watthour=518.4kwh"

1 unit =10rupees

Total cost

"Ru=518.4\times10=5184rupees"

Bulb replacement 15watt

"E=15\times4 watthour"

"E=6\times15\times4=360 watt hour"

Month

"E=360\times30 =10800watthour"

Annual


"E=10800\times12=12960 0watthour=12.96kwh"

Total cost


"Ru=129.6\times10=1296ruppes"

Saving(ru)=

"5184-1296=3888ru"